# SLayR: Scene Layout Generation with Rectified Flow

### v1

@@ Abstract
We introduce SLayR, Scene Layout Generation with Rectified flow. State-of-the-art text-to-image models achieve impressive results. However, they generate images end-to-end, exposing no fine-grained control over the process. SLayR presents a novel transformer-based rectified flow model for layout generation over a token space that can be decoded into bounding boxes and corresponding labels, which can then be transformed into images using existing models. We show that established metrics for generated images are inconclusive for evaluating their underlying scene layout, and introduce a new benchmark suite, including a carefully designed repeatable human-evaluation procedure that assesses the plausibility and variety of generated layouts. In contrast to previous works, which perform well in either high variety or plausibility, we show that our approach performs well on both of these axes at the same time. It is also at least 5x times smaller in the number of parameters and $37\%$ faster than the baselines. Our complete text-to-image pipeline demonstrates the added benefits of an interpretable and editable intermediate representation.
Figure 1: We introduce SLayR, a method for scene layout generation via rectified flow. SLayR generates editable and interpretable scene layouts, which can be rendered using conditional image diffusion. We demonstrate high plausibility and high variety of our generations.
@@ 1 Introduction
Recent advances in text-to-image modeling have focused on training denoising diffusion models to generate images from a prompt encoding and image noise [34, 35, 36, 6, 50]. While these methods provide very impressive results, we demonstrate that they show a lack in the variety of generated scenes when compared with images generated with scene layout conditioning. Furthermore, they lack the interpretability and controllability that a two stage text-to-layout followed by layout-to-image pipeline would natively provide. At the same time, many works have focused on generating images from given layouts [45, 51, 27] but these methods cannot perform general text-to-image generation without a given input layout.
Therefore, we introduce a novel text-to-layout generation model and combine it with an existing conditional image generation model as shown in Figure 1 to form a complete text-to-image pipeline, creating more varied images than the end-to-end image generation baseline with the benefit of a human-interpretable and controllable intermediate representation. Given a CLIP [33] embedding of a global scene prompt, we generate the layout using rectified flow [24], with a modified Diffusion Transformer (DiT) [32]. As baselines, we apply LayoutTransformer [11] with a small adaptation to handle global conditioning, and GPT4o  [29], as large language models (LLMs) also develop abstract representations of scenes and can be used to generate layouts [40, 21, 8]. However, we do not consider LLMs as a direct baseline, as they are trained on different and substantially larger datasets. Finally, we apply InstanceDiffusion [45] to obtain images.
We evaluate the final generated images using traditional metrics, and observe that we achieve state-of-the-art in FID [12] and KID [3] among text-to-layout generation baselines. By performing additional tests on images generated with no or fixed layouts, we find that interestingly, the CMMD [15] and IS [37] metrics indicate high quality on this limited data and thus cannot be considered conclusive for evaluating images that contain multi-object scenes.
@@
To establish a comparison directly on the quality of generated layouts, we introduce novel metrics to measure their plausibility and variety. Additionally, to achieve a reliable assessment, we introduce a repeatable human-evaluation study and use it to validate our findings. We show that our model yields the highest trade-off in generating layouts that are both, diverse and plausible, while being significantly more light-weight and faster than its competitors. Finally, we show that our approach can be conditioned with partial layouts and the native editing capabilities of the two-stage pipeline. In summary, our contributions are as follows:
@@
  • We propose the first model for text-to-layout generation using rectified flow, which can generate complex scene layouts with varying number and types of objects from a generic text prompt, and supports partial layout conditioning.
@@
  • We establish a suite of numerical metrics for generated layout evaluation and a well-designed human-evaluation study that can be used to assess the quality and be repeated by others.
  • We demonstrate that our intermediate layouts result in substantially improved variety of the generated scenes, while having high plausibility at the same time.
  • We demonstrate that our method provides controllability, and that the resulting images can be edited.
@@ 2 Related Work
Controlled Text-to-Image Generation. Denoising diffusion modeling initiated a new wave [42, 13, 43] of state-of-the-art text-to-image models  [35, 38]. In practice, it is desirable to have fine-grained user control over generated images that surpasses the capabilities of only conditioning on a single text prompt. One way of achieving this is to structure the latent generation [14, 18, 31], allowing inter- and extrapolation of latents. Our method follows an alternative paradigm of conditional generation [51, 27, 41, 28, 47]. In particular, we consider a conditioning on both a global and local bounding box prompts, following previous works GLIGEN [20] and Instance Diffusion [45]. The input to these methods has to be specified manually though, and therefore our work focuses on actually obtaining these layouts from a simple text prompt. We then leverage the existing InstanceDiffusion [45] approach to transform the generated layout to an image.
Layout Generation. LayoutTransformer  [11] is an autoregressive transformer which completes a partial sequence of object bounding boxes to form an image layout. We argue that a model that generates the layout jointly is a more natural choice, as the task does not provide an intrinsic ordering.
A work that leverages diffusion for layout generation is LayoutDM [4], but this model does not support global prompt conditioning, requiring specification and prediction of bounding boxes for all objects in the scene. LayoutFlow [10] is also related to our work as it generates User Interface (UI) layouts via Flow Matching [23]. However, this model is specialized to the UI layout domain, and thus does not natively support global conditioning. The same applies to other UI layout generators, such as DLT  [19] and LayoutFormer++  [16]. SceneGenie  [7] utilizes SG2Im  [17] to generate a scene layout from a pre-existing scene graph. As we assume only a simple global prompt, it is not applicable to our setting.
LLMs in Layout Generation. LLMs have been shown to address related layout generation tasks. LLM-grounded Diffusion [21] takes a multi-object prompt and generates a scene layout, with a bounding box for each object in the prompt. LayoutGPT [8] addresses this task as well, but using a different prompt template. Other works have demonstrated that LLMs can also reason about creating 3D layouts [53, 9, 49, 54, 1]. However, we do not focus on 3D scene generation and consider these works out of scope.
To reasonably measure an LLM’s aptitude for our task, we prompt GPT4o using our own template written to encourage chain of thought reasoning [46]. We measure against GPT4o’s performance in the layout generation task, although it was trained on significantly more data and cannot be considered a direct baseline.
Figure 2: Method Overview. Our layout generation model takes a set of noisy tokens and a global PCA-reduced CLIP embedding as input and uses a transformer-based rectified flow architecture to denoise the tokens into a scene layout that can then be transformed into images with InstanceDiffusion [45]. In order to decode the PCA-reduced CLIP embeddings back to be interpreted by humans, we compare them with a bank of predefined labels. The explicit intermediate layout representation of our pipeline allows editing the scene.
@@
Rectified Flow. Diffusion modeling has inspired numerous variants and improvements, one of which is rectified flow [24]. Prior works on the related text-to-image generation task [25, 6] have successfully leveraged it in their pipelines, with the added benefit of not requiring a noise schedule like DDIM [43] or DDPM [13]-based approaches. This motivates us to use rectified flow for our model.
@@ 3 Method
The central part of our work is the text-to-layout generation module, which we combine with the existing InstanceDiffusion [45] to form a complete text-to-image pipeline. An overview is provided in Fig. 2.
@@ 3.1 Rectified Flow Preliminaries
The model $v$ is trained to take $(\mathbf{x}_{i}(t),t)$ as input and predict the derivative of the path between $\mathbf{x}_{i}(0)$ and $\mathbf{x}_{i}(1)$, which according to Equation 1 is $\mathbf{x}_{i}(1)-\mathbf{x}_{i}(0)$. The model is trained with the following objective:
@@
$\mathbf{x}_{i}(1)=\mathbf{x}_{i}(0)+\sum_{t=1}^{T}v(x((t-1)/{T}),t/T)\cdot\frac{1}{T}\,.$ (3)
@@ 3.2 Layout Representation
We start with defining a scene representation as the basis for our generative architecture. A training sample $(\mathbf{x}_{i},P_{i})$ is composed of a global image caption prompt $P_{i}$ and a set of $J$ object tokens $\mathbf{x}_{i}=\{\mathbf{x}_{i}^{j}\in\mathbb{R}^{d+5}\}_{j\in J}$. The token representation of any single object in the scene is composed of
@@ 3.3 Our Model Architecture
Using Eq. 3 to sum over all timesteps yields the final layout $\{\mathbf{x}_{i}^{j}(1)\}_{j\in J}$ with PCA-reduced CLIP embeddings, bounding boxes, and opacities. Tokens with $\alpha_{i}^{j}(1)<0.5$ are considered unused and discarded. To determine object labels from the produced tokens, we unproject each $\mathbf{c}_{i}^{j}(1)$ from the PCA space into the CLIP feature space, and find the nearest CLIP embedding from a bank to produce interpretable labels. For image generation, we pass the unprojected embeddings directly into the downstream image generation module.
We explain how we adapt the RePaint [26, 39] technique for rectified flow to enable partial layout conditioning in the supplemental material. As we show in Sec. 5.8, our model can accommodate partial layouts where only some boxes or some labels are given.
@@ 3.4 Training
To construct a training sample from the ground-truth image layout $i$, we create ${\mathbf{c}_{i}^{j}}$ and ${\mathbf{b}_{i}^{j}}$ for each bounding box $j$, and initialize $\alpha_{i}^{j}:=1$. To ensure a consistent amount of tokens, we pad the samples by adding tokens with $\alpha_{i}^{j}:=0$, and ${\mathbf{c}_{i}^{j}}$ and ${\mathbf{b}_{i}^{j}}$ sampled from Gaussian noise.
@@
4 Metrics
Assessing the quality of our generated layouts is challenging. Solely evaluating the final images is not meaningful, as their quality is heavily impacted by the choice of conditional image generator and in addition we reveal in Sec. 5.5 that existing metrics are not conclusive. In the following, we therefore contribute novel suitable metrics that allow a reliable evaluation.
4.1 Generated Layout Numerical Metrics
To assess the quality of the generated layout distribution, we introduce four metrics that compare the plausibility and the variety of generated objects.
@@ 4.1 Generated Layout Numerical Metrics
We introduce $l_{\text{Pos}}^{(1)}$, and $l_{\text{Pos}}^{(2)}$ to measure how plausible the objects in a generated layout are arranged. Let $m$ index all bounding boxes of object-label $\ell$ for prompt $i$. For each object-label $\ell$, we obtain a distribution $k_{i}^{\ell}$ with KDE of the object’s bounding box $({\mathbf{b}_{i}^{\ell}})_{m}$ in all layouts with global prompt $i$. We compute the average likelihood over all objects and all global prompts, to measure the first order positional likelihood:
@@
We also want to measure the likelihood of spatial relationships between objects. Let $m^{*}$ index all bounding boxes of object-label ${\ell}^{\prime}$ For each object-labels pair $(\ell,{\ell}^{\prime})$, we obtain a distribution estimated with KDE $k_{i}^{{\ell},{{\ell}^{\prime}}}$ for the difference in the bounding box dimensions. We compute the average likelihood over all objects and all global prompts from our distributions to measure the second order positional likelihood:
@@
If this metric is small, this means that the variance is low.
4.2 Human Evaluation
Previous metrics capture whether objects appear in similar frequencies, positions, and stand in pairwise relationships similar to the ground-truth distribution. Truly assessing a generated layout requires full spatial and semantic interpretation that can only be carried out by humans. To this end, we introduce a human-evaluation study to assess the plausibility and variety of the final generated images.
To conduct the evaluation, we survey $60$ participants by presenting them images from different methods. For each method, each participant rates $10$ collections for their plausibility, and different $10$ collections for their variety, on a Likert scale ($1$ to $5$, where $1$ corresponds to very implausible/very low variance, and $5$ corresponds to very plausible/very high variance). For plausibility, we instructed participants to consider the overall realism of the collection, as well as how effectively it depicts the global text prompt. For variety, we instructed users to consider the spatial arrangement of objects in an image and implied camera angle in addition to overall image appearance.
A single evaluation of the study on Connect CloudResearch costs about \$160 and can be repeated by others. We provide the full details of the study in the supplemental material. The study was approved by the Ethics Review Board of our institution, ensuring compliance with ethical standards.

### v2

@@ Abstract
We introduce SLayR, Scene Layout Generation with Rectified flow, a novel transformer-based model for text-to-layout generation which can then be paired with existing layout-to-image models to produce images. SLayR addresses a domain in which current text-to-image pipelines struggle: generating scene layouts that are of significant variety and plausibility, when the given prompt is ambiguous and does not provide constraints on the scene. SLayR surpasses existing baselines including LLMs in unconstrained generation, and can generate layouts from an open caption set. To accurately evaluate the layout generation, we introduce a new benchmark suite, including numerical metrics and a carefully designed repeatable human-evaluation procedure that assesses the plausibility and variety of generated images. We show that our method sets a new state of the art for achieving both at the same time, while being at least 3$\times$ times smaller in the number of parameters.
Figure 1: Left: We introduce SLayR, a method for scene layout generation via rectified flow. Middle: SLayR generates interpretable and editable scene layouts, which can be rendered using a layout-to-image generator. Right: Our method sets a new state of the art in generating more varied and plausible scenes than other methods and LLMs.
@@ 1 Introduction
Recent advances in text-to-image modeling have focused on training denoising diffusion models  [43, 12, 44] to generate images from a prompt encoding and image noise [37, 38, 39, 6, 55, 40], as well as incorporating finer-grained control modalities [13, 18, 34, 57, 30, 42, 31, 49]. Building upon these advancements, prior works have demonstrated the editability and interpretability advantages of a multistage text-to-layout-to-image model, where the user can view and manipulate an intermediate layout consisting of bounding boxes for object-level scene elements [23, 7, 60, 8, 54, 61, 1]. These works use LLMs as text-to-layout generators, and focus on parsing multi-object prompts (e.g. two dogs next to a cat). However, a closer inspection reveals that these models do not generate high variety (see Figure 1, right) or collapse entirely (see Figure 2), when presented with prompts that have few constraints and leave a high degree of ambiguity.
This motivates us to propose a novel light-weight text-to-layout generation model for expanding unconstrained prompts (e.g. a park, a beach) into a variety of plausible scene layouts (see Figure 1, left and middle). The model cannot rely on the prompt to deliver guiding information, and must instead have extensive knowledge about the general distribution of scene structures. Given a CLIP [36] embedding of a global scene prompt, we generate the layout using rectified flow [27], with a Diffusion Transformer (DiT) [35]. We test our model’s performance across a set of layout-to-image generation models  [46, 22, 50, 23]. Since the domain of unconstrained text-to-layout generation is largely unexamined for non-LLM approaches, we also evaluate a comprehensive set of baseline architectures adapted from UI/document generation.
Ours LayoutGPT Ours LayoutGPT
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Degenerated layouts from unconstrained prompts of a skyscraper, and a street from LayoutGPT [7] vs. our layouts. LayoutGPT shows to produce degenerate scenes 2/3rds of the times while our produces detailed plausible scene layouts.
@@
We pair SLayR with several available layout-to-image model generation models and achieve optimal scores in CMMD [15], FID [11], KID [3], VQA  [26], HPSv2  [48], and ImageReward  [52]. As these metrics evaluate only the quality of the images and not the layouts, we furthermore introduce a set of novel metrics to statistically assess the plausibility and variety of scene layouts. To validate our findings, we introduce a comprehensive, repeatable human-evaluation study. We show that, across all examined text-to-layout-to-image pipelines, our model yields the best trade-off in generating images that are both diverse and plausible, while being significantly more light-weight than its competitors. Our approach can be conditioned on partial layouts and directional constraints, and enjoys the native editing capabilities of the two-stage pipeline. In summary, our contributions are:
@@
  • We propose the first general purpose model for text-to-layout generation using rectified flow, which can generate complex scene layouts with varying number and types of objects from a generic text prompt, supports partial layout conditioning, directional constraints, and can scale to open scene captions.
@@
  • We establish a comprehensive suite of numerical metrics for generated layout evaluation and a well-designed human-evaluation study that can be used to assess the layout quality and be repeated by others.
  • We demonstrate that our method can successfully integrate into a text-to-layout-to-image pipeline, yielding the state-of-the-art in achieving both variety and plausibility.
Upon publication, the source code and models will be made available to the community.
@@ 2 Related Work
LLMs in Scene Layout Generation. In 2D layout generation, LLM-Grounded Diffusion [23] and LayoutGPT [7] generate scene layouts from multi-object prompts, with a bounding box for each object in the prompt, using prompt templates that facilitate in-context learning. We found that directly plugging in single-word prompts into these templates yields degenerate solutions (layouts with a single bounding box labeled with the prompt) in approximately 2/3 of queries (see Figure 2). Therefore, to reasonably measure an LLM’s aptitude for our task and to encourage chain-of-thought reasoning [47] we adapt LLM-Grounded Diffusion’s template with in-context examples to our domain as a baseline. We note here also that GPT4o is a vastly larger model that was trained on much larger corpora of data than other baselines.
Adapting UI Generation. Our task of scene layout generation is distinct from User Interface (UI) generation: scene and object captions are from open sets, whereas UI layouts lack global captions and have labels from a small fixed set. Nevertheless, they can serve as interesting baselines, and we adapt several of these models using their conditional generation capabilities. We use LayoutTransformer  [10] as a representative for autoregressive transformer approaches, which completes a partial sequence of object bounding boxes to form an image layout. Works such as LayoutFormer++  [16] extend LayoutTransformer with added conditioning, but this is not the focus of our assessment of adapted UI generation, and these baselines are considered redundant. We also adapt LayoutDM  [14] and LayoutFlow  [9] as representative baselines for diffusion-based methods for UI generation [56, 4, 20]. For GAN-based approaches [21], while LayoutGAN++ [17] supports inter-bounding-box relationships, the Lagrange multiplier constraint formulation cannot be adapted to support global conditioning. In contrast to our method, UI generation models by design do not extend into the open world scenario.
Figure 3: Method Overview. Our layout generation model takes a set of noisy tokens and a prompt encoded as a global CLIP embedding as input. The tokens are partitioned into bounding box information ${\mathbf{b}_{i}^{j}}$, reduced clip embeddings ${\mathbf{c}_{i}^{j}}$ and opacities $\alpha_{i}^{j}$, that are subsequently denoised using a transformer. For visualization purposes, the user can query the generated layout with labels and edit boxes by adding, moving or removing them. Finally, the generated layout is passed through an off-the-shelf layout-to-image generator.
@@
Rectified Flow. Diffusion modeling has inspired numerous variants and improvements, one of which is rectified flow [27]. Prior works on the related text-to-image generation task [28, 6] have successfully leveraged it in their pipelines, with the added benefit of not requiring a noise schedule like DDIM [44] or DDPM [12]-based approaches, motivating us to use it for our task.
Layout-to-Image Generation. We demonstrate that SLayR integrates well into downstream conditional diffusion models to form a complete text-to-image pipeline, with the added benefits of an interpretable and controllable intermediate layout phase. To control for the effect which the image generator has on the final generated image, we evaluate our layouts across multiple layout-to-image models. Although there are a wide variety of such models, [5, 53, 58, 45, 2, 51] we select four which are publically available and have been used successfully with LLM-driven layouts [23, 7] or have shown SOTA performance: InstanceDiffusion  [46], GLIGEN [22], BoxDiff [50], and LMD+ [23].
@@ 3 Method
The central part of our work is the text-to-layout generation module, which we combine with the existing layout-to-image generators  [46, 22, 23, 50] to form a complete text-to-image pipeline. An overview is provided in Fig. 3.
@@ 3.1 Rectified Flow Preliminaries
The model $v$ is trained to take $(\mathbf{x}_{i}(t),t)$ as input and to predict the derivative of the path between $\mathbf{x}_{i}(0)$ and $\mathbf{x}_{i}(1)$, which according to Equation 1 is $\mathbf{x}_{i}(1)-\mathbf{x}_{i}(0)$. The model is trained with the following objective:
@@
$\mathbf{x}_{i}(1)=\mathbf{x}_{i}(0)+\sum_{t=1}^{T}v(\mathbf{x}_{i}(\frac{t-1}{T}),\frac{t}{T})\cdot\frac{1}{T}\,.$ (3)
@@ 3.2 Layout Representation
We start with defining a scene representation as the basis for our generative architecture. A training sample $(\mathbf{x}_{i},P_{i})$ is composed of a global image caption prompt $P_{i}$ and a set of $J$ object tokens $\mathbf{x}_{i}=\{\mathbf{x}_{i}^{j}\in\mathbb{R}^{d+5}\}_{j\in J}$. The token representation of any single object is composed of
@@ 3.3 Our Model Architecture
Using Eq. 3 to sum over all timesteps yields the final layout $\{\mathbf{x}_{i}^{j}(1)\}_{j\in J}$ with PCA-reduced CLIP embeddings, bounding boxes, and opacities. Tokens with $\alpha_{i}^{j}(1)<0.5$ are considered unused and discarded. For image generation, we unproject each $\mathbf{c}_{i}^{j}(1)$ from the PCA space into the CLIP feature space and pass the embeddings directly into the downstream image generation module. We decode labels into text for visualization purposes only, where the decoding can be achieved through label queries from the user or from a label bank.
3.4 Partial Conditioning and Constraints
We explain how we adapt the RePaint [29, 41] technique for rectified flow to enable partial layout conditioning in the supplemental material. As we show in Sec. 5.5, our model can accommodate partial layouts where only some boxes or labels are given. Additionally, we can impose inter-bounding box positional constraints by adding a directional drift on the bounding boxes during inference. We provide a brief sketch here and further details in the supplement. Let $\mathbf{y}_{i}$ be the partial layout, $M$ be a conditioning mask, and $\mathbf{d}_{i}$ be a vector specifying all drift conditions, all with the same dimension as $\mathbf{x}_{i}(t)$. During inference, we initialize $\mathbf{x}_{i}(0)$ from Gaussian noise, and at timestep $t$, we first apply the rectifed flow inference step:
$\mathbf{x}_{i}(\frac{t}{T})\leftarrow\mathbf{x}_{i}(\frac{t-1}{T})+v(\mathbf{x}_{i}(\frac{t-1}{T}),\frac{t}{T},P_{i})\cdot\frac{1}{T}\,,$ (5)
then compute time-adjusted partial conditioning:
$\mathbf{y}_{i}(\frac{t}{T})\leftarrow\mathbf{y}_{i}\cdot\frac{t}{T}+\mathbf{x}_{i}(0)\cdot\frac{1-t}{T}\,,$ (6)
and finally apply conditioning:
$\mathbf{x}_{i}(\frac{t}{T})\leftarrow\mathbf{x}_{i}(\frac{t}{T})\odot(1-M)+\mathbf{y}_{i}(\frac{t}{T})\odot M+\mathbf{d}_{i}\,.$ (7)
In addition to handling unconstrained or explicitly constrained prompts, following prior work [7], we integrate the ability to extract constraints from complex text prompts using LLMs by leveraging the LMM to extract the objects and positional constraints from them. We show that this integration allows us to perform well in spatial and numerical reasoning tasks in Sec. 5.4.
@@ 3.5 Training
To construct a training sample from the ground-truth image layout $i$, we create ${\mathbf{c}_{i}^{j}}$ and ${\mathbf{b}_{i}^{j}}$ for each bounding box $j$, and initialize $\alpha_{i}^{j}:=1$. To ensure a consistent amount of tokens, we pad the samples by adding tokens with $\alpha_{i}^{j}{\mathbf{b}_{i}^{j}}:=0$, and ${\mathbf{c}_{i}^{j}}$ to the null string embedding.
@@
4 Novel Metrics
4.1 Human Evaluation of Generated Images
To compliment generated image metrics  [15, 11, 3, 26, 48, 52] which we measure in  Sec. 5.2, we argue that the true plausibility and variety of images can only be assessed by humans and therefore introduce a human-evaluation study.
To conduct the evaluation, we survey $60$ participants by presenting them images from text-to-layout-to-image models with different text-to-layout baselines. For each method, each participant rates $10$ collections for their plausibility, and different $10$ collections for their variety, on a Likert scale ($1$ to $5$, where $1$ corresponds to very implausible/very low variance, and $5$ corresponds to very plausible/very high variance). For plausibility, we instructed participants to consider the overall realism of the collection, as well as how effectively it depicts the global text prompt. For variety, we instructed users to consider the spatial arrangement of objects in an image and implied camera angle in addition to overall image appearance. We control for the choice of layout-to-image generator by running our evaluation once for each of our tested layout-to-image generators (InstanceDiffusion  [46], GLIGEN [22], BoxDiff [50], and LMD+ [23]). We provide the full details of the study in the supplemental material. The study was approved by the Ethics Review Board of our institution, ensuring compliance with ethical standards.
4.2 Generated Layout Metrics
In the following, we introduce four metrics to assess the generated scenes layouts’ plausibility and variety.
@@ 4.2 Generated Layout Metrics
We introduce $l_{\text{Pos}}^{(1)}$, and $l_{\text{Pos}}^{(2)}$ to measure how plausible the objects in a generated layout are arranged. Let $m$ index all bounding boxes of object-label $\ell$ for prompt $i$. For each object-label $\ell$, we obtain a distribution $k_{i}^{\ell}$ with KDE of the object’s bounding box $({\mathbf{b}_{i}^{\ell}})_{m}$ in all layouts with global prompt $i$. We compute the average likelihood over all objects and all global prompts, to measure the first-order positional likelihood:
@@
We also want to measure the likelihood of spatial relationships between objects. Let $m^{*}$ index all bounding boxes of object-label ${\ell}^{\prime}$. For each object-label pair $(\ell,{\ell}^{\prime})$, we obtain a distribution estimated with KDE $k_{i}^{{\ell},{{\ell}^{\prime}}}$ for the difference in the bounding box dimensions. We compute the average likelihood over all objects and all global prompts from our distributions to measure the second order positional likelihood:
@@
If this metric is small, it means that the variance is low.
Adapting UI Generation Metrics.
Typical UI generation metrics must be reconsidered with shift in domain. Alignment [19] and Overlap [21] scores are not salient, as real world images often have misaligned or overlapping bounding boxes. The layout-FID [11] metric requires a GAN discriminator to compute, which we do not have in this new domain. We compute a standard max IoU (mIoU) [17] averaged across sampled scene categories.